OpenStudy (anonymous):

Erin bought 3 gallons of juice and 7 pounds of carrots and spent 4.80. jerry bought 5 gallons of juice and 6 pounds of carrots and spent 6.30. Represent the information as a pair of linear equations and determine the cost of 1 pound of carrots and the cost of a gallon of juice. please help!!!!!!!!

OpenStudy (anonymous):

First figure out what you need to find out.

OpenStudy (anonymous):

You need to know a) the cost of 1 pound of carrots and b) the cost of a gallon of juice.

OpenStudy (anonymous):

So, what you don't know should represented as the variables. Make one y and one x.

OpenStudy (anonymous):

y=3x(7)4.80?

OpenStudy (anonymous):

Okay, why don't we think of this in words first? Let's say that the cost of a gallon of juice is x, and the cost of a pound of carrots is y.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now Erin bought 3 gallons and 7 pounds, bringing his total cost up to 3x and 7y, correct? Since we know that this total is equal to 4.80, that means that 3x + 7y = 4.80. Do you understand why?

OpenStudy (anonymous):

yes because the carrots and juice is representative of the 3 and the 7

OpenStudy (anonymous):

Right. So knowing that, can you make an equation for Jerry's purchase now?

OpenStudy (anonymous):

well jerry's would be the 5x-6=6.30

OpenStudy (anonymous):

I think you made a typo there.

OpenStudy (anonymous):

haha yeah i did

OpenStudy (anonymous):

wait so jerry bought 5 gallons of juice and 6 pounds of carrots and spent 6.30

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so 5x+6y=6.30

OpenStudy (anonymous):

Exactly. Now, you can apply the same concept as the last question to isolate y, then combine the equations to solve for x.

OpenStudy (anonymous):

ohh ok